# Image Can Bring Your Memory Back: A Novel Multi-Modal Guided Attack against Image Generation Model Unlearning

### v1

@@ 3.4. Problem Formulation
We introduce a new attack strategy, which optimizes the image prompt, leveraging the multi-modal guidance (natively supported by the Stable Diffusion (Rombach et al., 2022)) to bypass unlearning mechanisms and regenerate the forgotten content.
@@
Specifically, given an unlearned image generation model (IGM) $\mathcal{G}_{u}$ that has been updated to suppress content associated with target concept $c$, a text prompt $P_{text}$ containing $c$ and an image $P_{image}$, our goal is to find an adversarially optimized image input $P_{image}^{adv}$ combined with a textual prompt $P_{text}$, which can trigger the unlearned IGM $\mathcal{G}_{u}$ to still generate content related to $c$. The output from the model is expressed as:
(7) $I^{*}=\mathcal{G}_{u}(P_{image}^{adv},P_{text}),\quad\text{s.t.}\quad I^{*}\approx I=\mathcal{G}(P_{text}),$
where $I^{*}$ maintains semantic similarity with image $I$, which is generated by the original model $\mathcal{G}$ with the text prompt $P_{text}$.
The adversarial optimization problem to obtain $P_{image}^{adv}$ can be formulated as:
(8) $P_{image}^{adv}=\arg\min_{P_{image}}\mathcal{L}_{adv}\left(\mathcal{G}_{u}(P_{image},P_{text}),\mathcal{G}(P_{text})\right).$
@@
Unlike prior attacks that modify the text prompt $P_{text}$, our method optimizes $P_{image}$ while keeping $P_{text}$ unchanged, ensuring that the attack does not compromise the semantic intent of the prompt. The optimization follows a gradient-based approach:
(9) $P_{image}^{adv}\leftarrow P_{image}-\eta\cdot\nabla_{P_{image}}\mathcal{L}_{adv}(\mathcal{G}_{u}(P_{image},P_{text}),\mathcal{G}(P_{text})),$
@@
where $\eta$ is the step size and $\mathcal{L}_{adv}$ is the adversarial loss. By solving this optimization problem, the adversarial image prompt $P_{image}^{adv}$ with the given text prompt $P_{text}$ can effectively exploit vulnerabilities in the unlearned model and restore the forgotten content while maintaining high semantic alignment with $P_{text}$.
Figure 2. Overview of the Recall framework. Stage I: the reference image $P_{ref}$ and an initialized image prompt $P_{image}^{init}$—constructed via noise blending—are encoded into latent representations $z_{ref}$ and $z_{adv}$, respectively, using the image encoder $\mathcal{E}_{i}$ of the unlearned model $\mathcal{G}_{u}$. Stage II: the diffusion process is simulated by passing both latents through the U-Net $\mathcal{F}_{\theta}$ across timesteps $t$, yielding predicted noise estimates $\hat{\epsilon}_{ref}$ and $\hat{\epsilon}_{adv}$. An adversarial loss $\mathcal{L}_{adv}$ is computed based on their discrepancy and used to iteratively update $z_{adv}$ in a PGD (Madry et al., 2018) manner. Stage III: the optimized latent $z_{adv}$ is decoded into an adversarial image $P_{image}^{adv}$, which is then paired with the given textual prompt $P_{text}$ and fed into the unlearned generative model $\mathcal{G}_{u}$. The final output image $I^{*}$ effectively reconstructs content associated with the erased concept, thereby revealing vulnerabilities in the unlearning mechanism under the multi-modal guidance.
4. Methodology
4.1. Overview
@@
We propose Recall, a novel multi-modal adversarial framework against unlearned image generation models (IGMs). Unlike conventional text-only attacks, Recall integrates adversarially optimized image prompts with the original textual inputs, leveraging a reference image $P_{ref}$—which implicitly contains the erased concept—as guidance throughout the optimization process. As illustrated in Figure 2, the framework consists of three stages. Stage I: Latent Encoding (Sec. 4.2). The reference image $P_{ref}$ and an initial image prompt $P_{image}^{init}$—constructed by injecting substantial random noise $\delta$ into $P_{ref}$—are encoded into latent representations $z_{ref}$ and $z_{adv}$ via the image encoder $\mathcal{E}_{i}$ in the unlearned IGM $\mathcal{G}_{u}$. Stage II: Iterative Latent Optimization (Sec. 4.3). The adversarial latent $z_{adv}$ is iteratively optimized under the guidance of the fixed reference latent $z_{ref}$. At each diffusion timestep $t$, the U-Net predicts noise residuals $\hat{\epsilon}_{ref}$ and $\hat{\epsilon}_{adv}$. The adversarial loss $\mathcal{L}_{adv}$—defined as the discrepancy between these predictions—is minimized through gradient-based updates. Stage III: Multi-modal Attack (Sec. 4.4). The optimized latent $z_{adv}$ is decoded into an adversarial image $P_{image}^{adv}$. When paired with the original text prompt $P_{text}$, this multi-modal input is fed into the unlearned IGM $\mathcal{G}_{u}$, resulting in a regenerated image $I^{*}$ that aligns visually with $P_{text}$ and successfully recovers the erased target concept $c$. Below, we elaborate on the design details of each stage. The pseudo-code of the Recall pipeline is shown in Alg. 1 of Appendix C.
4.2. Image Encoding
@@
To avoid incurring additional computational overhead from external classifiers or relying on the original IGM, we introduce a reference image $P_{ref}$ containing the target concept $c$—which can be sourced from the internet—to guide the generation process. This reference implicitly embeds the erased concept, thereby facilitating adversarial optimization of the initial image prompt $P_{image}^{init}$. To enhance efficiency and precision, Recall performs the optimization directly in the latent space representation $z_{adv}$ of the image prompt.
@@
As illustrated in Figure 2, we initialize $P_{image}^{init}$ by blending a small portion of the reference image $P_{ref}$ with random noise $\delta$ sampled from an isotropic Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$:
(10) $P_{image}^{init}\leftarrow\lambda\cdot P_{ref}+(1-\lambda)\cdot\delta,\quad\delta\sim\mathcal{N}(0,I),$
@@
To accelerate optimization, both $P_{image}^{init}$ and $P_{ref}$ are encoded into the latent space using the image encoder $\mathcal{E}_{i}$ from the unlearned model, yielding:
(11) $z_{i}=\mathcal{E}_{i}(P_{image}^{init}),\quad z_{ref}=\mathcal{E}_{i}(P_{ref}),$
@@
4.3. Iterative Latent Optimization
@@
4.3.1. Generation of Latent $z_{t}$.
Unlike standard latent diffusion, Recall generates the noisy latent at timestep $t$ as:
(12) $z_{t}=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$
@@
4.3.2. Optimization under Multi-Modal Guidance.
@@
(13) $\hat{\epsilon}_{ref}=\mathcal{F}_{\theta}(z_{\{ref,t\}},t,h_{t});\quad\hat{\epsilon}_{adv}=\mathcal{F}_{\theta}(z_{\{adv,t\}},t,h_{t}).$
The discrepancy between these two noise predictions forms the basis of the adversarial objective function.
As discussed previously, our attack explicitly targets the latent representation $z_{adv}$ of the adversarial image prompt $P_{image}^{adv}$, aiming to efficiently induce the unlearned IGM model to regenerate the previously unlearned content. Specifically, at each diffusion timestep $t$, we iteratively refine the adversarial latent representation $z_{adv}$ using a gradient-based optimization procedure guided by the adversarial loss $\mathcal{L}_{adv}$. To enhance stability and facilitate convergence, we incorporate momentum-based gradient normalization into our optimization scheme (Dong et al., 2018). Specifically, we iteratively update the latent adversarial variable $z_{adv}$ over $N$ epoches according to:
(14) $v_{i}=\beta\cdot v_{i-1}+\frac{\nabla_{z_{adv}}\mathcal{L}_{adv}}{\|\nabla_{z_{adv}}\mathcal{L}_{adv}\|_{1}+\omega},\quad z_{adv}\leftarrow z_{adv}+\eta\cdot\text{sign}(v_{i}),$
@@
(15) $z_{adv}\leftarrow z_{adv}+\gamma\cdot z_{ref},$
@@
4.3.3. Objective Function.
The adversarial objective function $\mathcal{L}_{adv}$ explicitly quantifies the discrepancy between noise predictions generated from the adversarial latent $\hat{\epsilon}_{adv}$ and reference latent $\hat{\epsilon}_{ref}$ with U-Net as step $t$, respectively:
(16) $\mathcal{L}_{adv}=\mathcal{M}(\hat{\epsilon}_{\{ref,t\}},\hat{\epsilon}_{\{adv,t\}})=\|\hat{\epsilon}_{\{ref,t\}}-\hat{\epsilon}_{\{adv,t\}}\|_{2}^{2},$
where $\mathcal{M}$ denotes a similarity measurement. In this work, we specifically employ the mean squared error (MSE).
4.3.4. Adversarial Image Reconstruction.
After optimization, the refined adversarial latent $z_{adv}$ is subsequently decoded into the image space through the image decoder $\mathcal{D}_{i}$ of the unlearned SD model to generate the final adversarial image used for the attack:
(17) $P_{image}^{adv}=\mathcal{D}_{i}(z_{adv}).$
4.4. Multi-modal Attack
Once the adversarial image $P_{image}^{adv}$ is obtained, we leverage the multi-modal conditioning mechanism of the unlearned model $\mathcal{G}_{u}$ to generate images containing the forgotten content and semantically aligned with the text prompt $P_{text}$. The final image generation process integrates both the optimized adversarial image prompt and the original text prompt in a multi-modal manner:
(18) $I^{*}=\mathcal{G}_{u}(P_{image}^{adv},P_{text}),$
@@ 4.4. Multi-modal Attack
5. Experiments
To thoroughly evaluate the effectiveness of our proposed Recall, we conduct extensive experiments involving TEN state-of-the-art unlearning techniques across four representative unlearning tasks: Nudity, Van Gogh-style, Object-Church, and Object-Parachute. These settings yield a total of forty unlearned IGMs based on Stable Diffusion v1.4. Our objective is to systematically validate the effectiveness and generalization of our proposed multi-modal guided attack against different scenarios.
Table 1. Attack performance of various attack methods against unlearned IGMs in four representative unlearning tasks, evaluated by ASR (%) and Avg. ASR (%). The best attack performance is highlighted in bold, while the second-best is underlined.
Task Method ESD FMN SPM AdvUnlearn MACE RECE DoCo UCE Receler ConcptPrune Avg. ASR
Nudity Text-only 10.56 66.90 32.39 1.41 3.52 7.04 30.99 8.45 8.45 73.24 24.30
Image-only 0 18.31 12.68 4.23 5.63 14.08 3.52 11.97 6.34 13.38 9.01

### v2

@@ 3.3 Problem Formulation
We introduce a new attack strategy that optimizes image prompts by leveraging multi-modal guidance, which is natively supported by Stable Diffusion (Rombach et al., 2022a), to bypass unlearning mechanisms and regenerate erased content.
@@
Given an unlearned image generation model (IGM) $\mathcal{G}_{u}$ that has been updated to suppress content associated with target concept $c$, a text prompt $P_{text}$ containing $c$, and an image $P_{img}$ relevant to the concept $c$, we aim to find an adversarial image input $P_{img}^{adv}$ such that, when paired with $P_{text}$, the unlearned IGM $\mathcal{G}_{u}$ generates image $I^{*}$ related to $c$:
$I^{*}=\mathcal{G}_{u}(P_{img}^{adv},P_{text}),\quad\text{s.t.}\quad I^{*}\approx I\mid c,$ (1)
where $I\mid c$ denotes images that explicitly contains the target concept $c$; these images can come from the original model $\mathcal{G}$ with the same text prompt $P_{\text{text}}$ or from any other source.
The adversarial image prompt $P_{img}^{adv}$ is obtained by solving:
$P_{img}^{adv}=\arg\min_{P_{img}}\mathcal{L}_{adv}\left(\mathcal{G}_{u}(P_{img},P_{text}),I\right),$ (2)
where $\mathcal{L}_{adv}$ is an adversarial loss function.
@@
Unlike prior attacks that modify the text prompt $P_{text}$, we optimize $P_{img}$ while keeping $P_{text}$ unchanged, thus preserving the semantic intent. The optimization follows a gradient-based approach:
$P_{img}^{adv}\leftarrow P_{img}-\eta\cdot\nabla_{P_{img}}\mathcal{L}_{adv}(\mathcal{G}_{u}(P_{img},P_{text}),I),$ (3)
@@
where $\eta$ is the step size. This process enables the adversarial image prompt $P_{img}^{adv}$, together with $P_{text}$, to exploit vulnerabilities in the unlearned model and recover the erased content while maintaining semantic alignment with the text prompt.
4 Methodology
4.1 Overview
@@
We propose Recall, a multi-modal adversarial framework targeting unlearned IGMs. Unlike conventional text-only attacks, Recall jointly optimizes adversarial image prompt by leveraging a reference image $P_{ref}$ as guidance. As illustrated in Figure 2, Recall comprises three stages: (1) Latent Encoding: The reference image $P_{ref}$ and a noise-injected initial prompt are encoded into latent representations. (2) Iterative Latent Optimization: The adversarial latent is iteratively refined under the guidance of the reference latent by minimizing the discrepancy between their predicted noise residuals. (3) Multi-modal Attack: The optimized latent is decoded to an adversarial image, which, paired with the text prompt, forms a multi-modal input to the unlearned IGM, enabling effective recovery of the erased target concept. Details of each stage are described in the following sections, and the overall pipeline is summarized in Algorithm 1 (Appendix B).
Figure 2: Overview of Recall. Given a reference image $P_{ref}$ that depicts the erased concept and a heavily noised initial image prompt $P_{img}^{init}$, we iteratively optimize the latent $z_{adv}$ (initialized from $P_{img}^{init}$) to align with the reference latent $z_{ref}$ under the same text condition. After optimization, $z_{adv}$ is decoded into an adversarial image $P_{img}^{adv}$, which is then paired with the original text prompt and fed into the unlearned model, enabling recovery of the erased concept, thereby exposing vulnerabilities of current unlearning mechanisms under multi-modal guidance.
4.2 Image Encoding
@@
To avoid incurring additional computational overhead from external classifiers or relying on the original IGM, we introduce a reference image $P_{ref}$ containing the target concept $c$ to guide the generation process, where the reference image $P_{ref}$ can be sourced from multiple sources, such the internet, public datasets, or self-collected. This reference implicitly embeds the erased concept, thereby facilitating adversarial optimization of the initial image prompt $P_{img}^{init}$. To enhance efficiency and precision, Recall performs the optimization directly in the latent space representation $z_{adv}$ of the image prompt.
@@
As illustrated in Figure 2, we initialize $P_{img}^{init}$ by blending a small portion of the reference image $P_{ref}$ with random noise $\delta$ sampled from an isotropic Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$:
$P_{img}^{init}\leftarrow\lambda\cdot P_{ref}+(1-\lambda)\cdot\delta,\quad\delta\sim\mathcal{N}(0,I),$ (4)
@@
To accelerate optimization, both $P_{img}^{init}$ and $P_{ref}$ are encoded into the latent space using the image encoder $\mathcal{E}_{i}$ from the unlearned model, yielding:
$z_{i}=\mathcal{E}_{i}(P_{img}^{init}),\quad z_{ref}=\mathcal{E}_{i}(P_{ref}),$ (5)
@@
4.3 Iterative Latent Optimization
@@
Generation of Latent $z_{t}$.
Unlike standard latent diffusion, which typically initializes from a randomly sampled latent, Recall generates the noisy latent at timestep $t$ as:
$z_{t}=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (6)
@@
Optimization under Multi-Modal Guidance.
@@
$\hat{\epsilon}_{ref}=\mathcal{F}_{\theta}(z_{\{ref,t\}},t,h_{t});\ \hat{\epsilon}_{adv}=\mathcal{F}_{\theta}(z_{\{adv,t\}},t,h_{t}).$ (7)
The discrepancy between two noise predictions forms the basis of the adversarial objective function.
As discussed previously, our attack explicitly targets the latent representation $z_{adv}$ of the adversarial image prompt $P_{img}^{adv}$, aiming to efficiently induce the unlearned IGM model to regenerate the previously unlearned content. Specifically, at each diffusion timestep $t$, we iteratively refine the adversarial latent representation $z_{adv}$ using a gradient-based optimization procedure guided by the adversarial loss $\mathcal{L}_{adv}$. To enhance stability and facilitate convergence, we incorporate momentum-based gradient normalization into our optimization scheme (Dong et al., 2018). Specifically, we iteratively update the latent adversarial variable $z_{adv}$ over $N$ epochs according to:
$v_{i}=\beta\cdot v_{i-1}+\frac{\nabla_{z_{adv}}\mathcal{L}_{adv}}{\|\nabla_{z_{adv}}\mathcal{L}_{adv}\|_{1}+\omega},\ z_{adv}\leftarrow z_{adv}+\eta\cdot\text{sign}(v_{i}),$ (8)
@@
$z_{adv}\leftarrow(1-\gamma)z_{adv}+\gamma\cdot z_{ref},$ (9)
@@
Objective Function.
The adversarial objective function $\mathcal{L}_{adv}$ explicitly quantifies the discrepancy between noise predictions generated from the adversarial latent $\hat{\epsilon}_{adv}$ and reference latent $\hat{\epsilon}_{ref}$ with U-Net at step $t$, respectively:
[…]
where $\mathcal{M}$ denotes a similarity measurement. In this work, we employ the mean squared error (MSE).
Adversarial Image Reconstruction.
After optimization, the refined adversarial latent $z_{adv}$ is subsequently decoded into the image space through the image decoder $\mathcal{D}_{i}$ of the unlearned SD model to generate the final adversarial image used for the attack: $P_{img}^{adv}=\mathcal{D}_{i}(z_{adv})$.
4.4 Multi-modal Attack
Once the adversarial image $P_{img}^{adv}$ is obtained, we leverage the multi-modal conditioning mechanism of the unlearned model $\mathcal{G}_{u}$ to generate images containing the forgotten content and semantically aligned with the text prompt $P_{text}$. The final image generation process integrates both the optimized adversarial image prompt and the original text prompt in a multi-modal manner:
$I^{*}=\mathcal{G}_{u}(P_{img}^{adv},P_{text}),$ (11)
@@ 4.4 Multi-modal Attack
Table 1: Attack comparisons against unlearned IGMs in six dataset for four representative unlearning tasks.
Task Method ESD FMN SPM AdvUnlearn MACE RECE DoCo UCE Receler ConceptPrune Avg. ASR
Nudity-I2P Text-only 10.56 66.90 32.39 1.41 3.52 7.04 30.99 8.45 8.45 73.24 24.30
Image-only 0.00 18.31 12.68 4.23 5.63 14.08 3.52 11.97 6.34 13.38 9.01
@@
Nudity-MMA Text-only 1.56 46.88 32.03 0.00 0.00 13.28 27.34 24.22 14.06 53.12 21.25
Text & R_noise 0.00 20.31 17.19 0.00 0.00 4.69 21.88 14.84 1.56 42.19 12.27
Text & Image 8.59 78.91 59.38 0.00 2.34 37.50 62.50 44.53 43.75 80.47 41.80
P4D-K 56.90 62.50 76.88 8.43 49.54 37.50 85.31 80.62 79.69 89.65 62.70
P4D-N 62.50 74.37 78.44 10.64 53.67 51.25 88.44 91.41 85.94 98.44 69.51
CCE 35.16 89.84 78.91 3.12 55.47 46.88 54.69 58.59 36.72 97.66 55.70
UnlearnDiffAtk 40.62 100.00 99.22 23.78 33.59 89.06 98.44 95.31 85.94 99.22 76.52
WACE-N 28.12 86.72 75.78 7.03 1.56 46.09 68.75 49.22 52.34 86.72 50.23
WACE-C 61.72 92.19 82.81 49.22 13.28 57.03 80.47 70.31 70.31 85.16 66.25
Recall 75.78 100.00 97.66 82.81 53.12 89.84 94.53 92.97 96.09 99.22 88.20
Nudity-ART Text-only 0.00 11.72 2.34 0.00 0.00 0.00 3.12 0.78 2.34 7.03 2.73
Text & R_noise 0.00 35.16 17.19 0.00 0.00 0.00 4.69 0.78 0.00 17.19 7.50
Text & Image 0.78 14.84 13.28 0.78 1.56 1.56 4.69 2.34 4.69 12.50 5.70
P4D-K 8.73 66.67 57.63 2.86 31.45 28.49 43.87 41.83 12.50 56.25 35.03
P4D-N 12.50 62.86 62.81 3.98 24.69 32.81 48.44 45.62 21.88 53.12 36.87
CCE 20.31 53.91 28.12 28.12 21.88 3.12 3.12 13.18 6.25 42.97 22.10
UnlearnDiffAtk 31.25 76.56 66.41 0.78 17.19 21.09 76.47 39.06 35.16 75.78 43.98
WACE-N 7.81 48.44 26.56 0.78 2.34 4.69 21.09 7.03 7.81 37.50 16.41
WACE-C 20.31 57.81 34.38 5.47 8.59 4.69 33.59 10.94 14.84 47.66 23.83
Recall 62.50 91.29 81.25 32.03 43.75 32.99 98.12 52.34 70.31 89.84 65.44
@@
P4D-K 6.00 40.00 24.00 2.00 4.00 14.00 72.00 18.00 20.00 96.00 29.60
P4D-N 36.00 82.00 70.00 8.00 22.00 12.00 52.00 14.00 2.00 84.00 38.20
CCE 74.00 92.00 72.00 48.00 54.00 34.00 52.00 52.00 38.00 88.00 60.40
WACE-N 30.00 84.00 46.00 10.00 10.00 6.00 26.00 4.00 6.00 88.00 31.00
WACE-C 56.00 84.00 60.00 6.00 16.00 8.00 32.00 22.00 14.00 90.00 38.80
5 Experiments
We conduct extensive experiments involving TEN SOTA unlearning techniques across four representative unlearning tasks: Nudity, Van Gogh-style, Object-Church, and Object-Parachute, thus yield a total of forty unlearned IGMs. Our objective is to systematically validate the effectiveness and generalization of our proposed multi-modal guided attack Recall against different scenarios.
5.1 Experimental Setup
Datasets. We evaluate on three nudity unlearning datasets (I2P (Schramowski et al., 2023), MMA (Yang et al., 2024), and ART (Li et al., 2024a)). For the remaining targets, such as Van Gogh-style, Object-Church, and Object-Parachute, we reuse the text prompts released by UnlearnDiffAtk to ensure protocol comparability. For methods that require a reference image, we provide one same additional image per unlearning task. Details of all prompts and reference images are provided in Appendix C.1, Table 3.
IGMU Methods. We evaluate our approach across ten state-of-the-art IGMU techniques: ESD (Gandikota et al., 2023), FMN (Zhang et al., 2024a), SPM (Lyu et al., 2024), AdvUnlearn (Zhang et al., 2024c), MACE (Lu et al., 2024), RECE (Gong et al., 2024), DoCo (Wu et al., 2025), Receler (Huang et al., 2024), ConceptPrune (Chavhan et al., 2025), and UCE (Gandikota et al., 2024). The weights of involved unlearned SD models are sourced from three primary origins: ① the AdvUnlearn GitHub repository [^1] https://github.com/OPTML-Group/AdvUnlearn, as described in (Zhang et al., 2024c); ② weights officially released by their respective authors, such as RECE (Gong et al., 2024), MACE (Lu et al., 2024) and DoCo (Wu et al., 2025); and ③ weights trained in-house using official implementations provided by ourselves.
Baselines. We compare our proposed Recall against several representative attack baselines: Text-only (text prompts only), Image-only (reference image prompt only), Text & R_noise (text with a noised image), Text & Image (text prompt and reference image), P4D (with two variants P4D-K and P4D-N) (Chin et al., 2024a), CCE (Pham et al., 2024), UnlearnDiffAtk (Zhang et al., 2024d), and WACE(with two variants WACE-N and WACE-C) (Lu et al., 2025). Their detailed descriptions and implementation can be found in Appendix C.2.
Evaluation Metrics. We assess the effectiveness of our attack using task-specific deep learning-based detectors and classifiers, including the NudeNet detector (Praneeth, 2023), a ViT-based style classifier (Zhang et al., 2024d), and an ImageNet-pretrained ResNet-50 (He et al., 2016). The primary metric is attack Success Rate (ASR, %) and average ASR for attack performance, average attack time (seconds, s) for computational efficiency, CLIP Score (Hessel et al., 2021) for quantifying semantic alignment between generated images and prompts, and LPIPS (Zhang et al., 2018), Inception Score (IS) (Salimans et al., 2016), and a DINO-based feature distance (Oquab et al., 2024) for generated image diversity. Throughout all tables, the best attack performance is highlighted in bold, while the second-best is indicated with underlining.